OpenStudy (anonymous):

Use the substitution method to solve.5x – y = 10y = 4x + 2

OpenStudy (precal):

5x-y=10 y=4x+2 sub 2nd eq into first

OpenStudy (precal):

5x-(4x+2)=10

OpenStudy (precal):

solve for x

OpenStudy (this):

y=7-2/17

OpenStudy (precal):

@this what is that you posted?